Suppose an additional horizontal force 11.0 N directed to the left is applied to mass m 2 in the system of masses in previous sample problem . Find the acceleration of the system and the tension in the rope . ( refer to illustrative problem ) m2 = 6.5kg ----T---- m1 = 4.5 kg ----F----
1
Expert's answer
2020-12-06T17:29:38-0500
"(m_1+m_2)a=F\\(6.5+4.5)a=11\\a=1\frac{m}{s^2}\\T=m_1a=(4.5)(1)=4.5\ N"